-32 + 3[12 ÷ (5 - 8)] can you help

Found 2 solutions by algebrahouse.com, unlockmath:
Answer by algebrahouse.com(1659)

-32 + 3[12 ÷ (5 - 8)]
= -32 + 3[12 ÷ (-3)] {subtracted 5 - 8}
= -32 + 3(-4) {divided 12 by -3}
= -32 - 12 {multiplied 3 by -4}
= -44 {subtracted}
www.algebrahouse.com

Answer by unlockmath(1688)

Hello,
When the order of operation is known then we can figure this out.
-32 + 3[12 ÷ (5 - 8)]
First do the parentheses:
-32 + 3[12 ÷ (-3)]
Then do the brackets:
-32 + 3[-4]
Multiply:
-32-12
We get:
-44
Make sense?
RJ
www.math-unlock.com